OpenStudy (anonymous):

Jeremiah is making a shake by mixing two different protein powders, measured in ounces. The strawberry-flavored powder has 4 grams of protein per ounce. The banana-flavored powder has 3 grams of protein per ounce. He wants the drink to have a total of 6 ounces of powder and contain 22 grams of protein.

OpenStudy (anonymous):

a. Create a system of linear equations to represent the situation that will determine the exact number of ounces needed for each power.

OpenStudy (anonymous):

b. Based on the system you wrote, which algebraic method of solving systems of equations will you choose to solve the system? Explain your choice.

OpenStudy (anonymous):

c. Solve the system algebraically, showing your work. Then interpret the solution by explaining what your solution represents in the context of the problem.

OpenStudy (anonymous):

@Ashleyisakitty

OpenStudy (anonymous):

First choose variables to represent ounces of each powder. e.g. S=ounces of strawberry; B=ounces of banana. You'll need two equations for two unknowns. One equation will represent the total number of ounces of powder, the other equation will represent the grams of protein.

OpenStudy (anonymous):

4x + 3x = 22?

OpenStudy (anonymous):

@CliffSedge

OpenStudy (anonymous):

Yes, that would be an equation representing grams of protein with x=strawberry and y=banana.

OpenStudy (anonymous):

So you need another equation to represent the total number of ounces of both powders.

OpenStudy (anonymous):

x + y = 6?

OpenStudy (anonymous):

That'll do it. There are now many different methods you could use to solve the system. I'd go with substitution, personally.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

is that it for part A?

OpenStudy (anonymous):

Yep, part A is just to write the system (the two equations).

OpenStudy (anonymous):

And Part B is substitution?

OpenStudy (anonymous):

@CliffSedge

OpenStudy (anonymous):

Part B is whatever you want, but you have to give a reason why you pick whatever method. I personally like substitution in this situation, but you might like something else.

OpenStudy (anonymous):

Maybe you'd rather solve by graphing or using Cramer's rule or Gauss-Jordan elimination (though that is a bit overkill on a 2by3 system . . .)

OpenStudy (anonymous):

could you explain how you would solve it with substitution?

OpenStudy (anonymous):

@CliffSedge

OpenStudy (anonymous):

Well, equation (2) x+y=6 is very simple, so would be easy to rewrite as y=(6-x). Then substitute that expression for y, (6-x) into equation (1) thusly: \[\large 4x+3y=22\rightarrow4x+3(6-x)=22.\] Solve for x, then go back to equation (2) to get y.

OpenStudy (anonymous):

ok thanks for all your help :D